I Consider the Taylor series for ex about a=1. al give its first three terms. bl Write the entire series in E- notation.
Answer #1

biven f(u) = eh en a-1 f(x) = pk {(n) = en (n)= e cuies (as first three terms of Taylor series : = f(a) + f(o) (x-a) +